# Data-Centric Foundation Models in Computational Healthcare: A Survey

### v1

@@ 4. Multi-Modal Data Fusion
Data fusion is a useful strategy to aggregating the information of various medical data modalities towards improved decision-making in healthcare. Standard fusion methods can be conceptually categorized into three types (Stahlschmidt et al., 2022; Huang et al., 2020): early fusion, joint fusion, and late fusion. In early fusion, data from different modalities are combined at the input and passed to the subsequent network (Park et al., 2020; Xie et al., 2019; Zhao et al., 2021; Suresh et al., 2017; Bichindaritz et al., 2021; Fu et al., 2020). Joint fusion processes data from each modality through independent networks before fusing their representative feature maps as the input of the subsequent network (Braman et al., 2021; Spasov et al., 2018; Kawahara et al., 2018). Late fusion processes data from each modality through individual networks and fuses their output to make the final outcomes (Reda et al., 2018; Qiu et al., 2018). However, conventional methods suffer from a lack of scalability, generalizability, and cross-modal understanding (Huang et al., 2020). For example, early and joint fusion methods often require sufficient task-specific training data and computational resources, while late fusion methods usually show weaknesses in multi-modal information integration (Budikova et al., 2017). Figure 4 illustrates the healthcare data modalities, data fusion strategies, and associated healthcare tasks.
@@ 4.1. Data Fusing via Joint-Modal Pre-Training
FMs can handle multiple modalities via pre-training on massive-scale paired multi-modal data in a joint-modal mode to obtain a high-level understanding of inter-modality relationships. For instance, CLIP (Radford et al., 2021), ALIGN (Jia et al., 2021), and Florence (Yuan et al., 2021) adopt image-caption pairs collected from the Internet. These models extract image and text features with separated networks and project both features into a shared latent space, enabling a joint training network. These models have shown great zero-shot transferability when testing on unseen image-text combinations, including medical images and healthcare reports (Liu et al., 2023f). In addition, GLIP (Li et al., 2022), CoCa (Yu et al., 2022), and FLAVA (Singh et al., 2022) introduce more advanced fusion modules, such as cross-attention and generative models, to better explore cross-modality interaction during pre-training. Besides vision-language models, there exist models for other modalities with joint-modal pre-training. For example, CLAP (Elizalde et al., 2023) is an audio-language model pre-trained on audio-caption pairs.
Joint-modal pre-training has been applied for healthcare FMs, especially in radiology, where chest X-ray images and reports are frequently paired. ConVIRT (Zhang et al., 2022) is pre-trained on numerous pairs of chest X-rays and radiology reports, fusing vision and language modalities via contrastive learning. BioViL (Boecking et al., 2022) trains a visual model together with an in-domain (CXR) language model by optimizing contrastive loss and masked language modeling. In addition, PubMedCLIP (Eslami et al., 2021), CheXzero (Tiu et al., 2022), and PLIP (Huang et al., 2023a) extend the CLIP model to image-text data in biomedicine, radiology, and pathology respectively. BiomedJourney (Gu et al., 2023) is a latent diffusion model (Rombach et al., 2022) trained on tuples of pre- and post-disease progression images from patients coupled with a GPT-generated natural language description of the progression, enabling the generation of counterfactual medical images.
@@ 4.2. Data Fusing via LLMs
Transformer-style LLMs possess powerful semantic understanding capability via the attention mechanism (Brown et al., 2020), which can be transferred to multi-modal settings. To be specific, data from different modalities can be aggregated as the prompt input of an LLM (i.e., a sequence of tokens). These combined multi-modal inputs are then fused through the Transformer blocks in the LLM, exchanging information via attention layers (Wang et al., 2023a; Liu et al., 2023a, b). For instance, Frozen (Tsimpoukelli et al., 2021) concatenates image features and text tokens as the input of a pre-trained generative language model. This design enables the vision encoder to update based on information from both modalities. Similarly, SimVLM (Wang et al., 2021) fuses image features and text tokens with BERT-like bi-directional attention layers and trains the whole model together. Flamingo (Alayrac et al., 2022) takes as input visual data interleaved with text and inserts gated xattn-dense layers into the language model to condition it on visual inputs. BLIP-2 (Li et al., 2023c) adopts a well-designed Q-Former to bridge the gap between pre-trained vision and language models before passing into the LLM. In addition, LLaVA (Liu et al., 2023b) feeds image features and language instructions into LLaMA to train a multi-modal chatbot. NExT-GPT (Wu et al., 2023a) is a unified agent understanding multiple modalities, including text, images, videos, and audio.
@@
Since LLMs show great potential in the medical field (Nori et al., 2023), these methods are applied to medical data fusion as well. Med-Flamingo (Moor et al., 2023b) is an open-ended MedVQA model based on Flamingo, pre-trained on paired and interleaved medical image-text data. MedVInT (Zhang et al., 2023d) applies instruction tuning to a LLaVA-like multi-modal model with MedVQA data to pre-train a generalizable MedVQA foundation model. Visual Med-Alpaca (Shu et al., 2023) is designed to handle multi-modal biomedical tasks, based on the LLaMa-7B architecture and trained with an instruction set designed cooperatively by GPT-3.5-Turbo and human experts. Moreover, LLaVA-Med (Li et al., 2023e) curates biomedical instruction-tuning data by prompting language-only GPT-4 to generate multi-round questions and answers about biomedical images. Surprisingly, GPT-4 generates high-quality visual question-answering conversations even if it has access only to the text. These models show outstanding zero-shot/few-shot performance on downstream medical tasks that require multi-modality understanding.
@@
Current multi-modal FMs predominantly focus on language and vision, given their ubiquitous use as the most common modalities. For FM applications in the healthcare field, there are only handful examples including genomics (Ding et al., 2023b), electrocardiogram (Chang et al., 2023a), electroencephalogram (Cui et al., 2023a), and tabular data (Wang et al., 2023c). We believe that joint-modal pre-training techniques, combined with LLMs’ logical reasoning capabilities, could benefit healthcare data fusion of modalities beyond the scope of language and vision.
@@ 5. Data Quantity
Towards applying FMs in the healthcare system safely and responsibly, we must address a variety of data quantity hurdles. The trade-off of data quantity is between the limited information provided by a constrained dataset and the essential information required for training a robust model on the specific healthcare task. Such a disparity can result in inadequate downstream model training, yielding non-robust, inaccurate, and even unreliable model outcomes. However, public patient data records are often scarce due to the rigorous patient privacy protection regulations. In addition, we must deal with the costly real-world dataset curation processes including data collection, cleaning, and annotation. While deep learning has shown its promise, the barrier of limited data access in healthcare continues to be a significant roadblock for scalable AI-based research.
@@ 6.1. Healthcare Text Annotation
Healthcare text annotation is a process that extracts and categorizes critical information within a wide range of healthcare texts. This process plays a crucial role in the healthcare system to enhance data quality and knowledge discovery (Xia and Yetisgen-Yildiz, 2012). However, healthcare text annotation is labor-intensive and time-consuming, observed from studies in medical condition labeling (Briesacher et al., 2008), patient record annotation (Hogan and Wagner, 1997), and key medical entity identification (Abacha and Zweigenbaum, 2011). To streamline a text annotation process, an emerging direction is to harness general-purpose LLMs. For instance, Gilson et al. (Gilson et al., 2022) reveal that ChatGPT (OpenAI, 2021) achieves comparable human-level performance on medical question-answering tasks, showcasing the potential for LLMs to enhance the efficiency of text annotation in a more flexible manner. Med-PaLM 2 (Singhal et al., 2023), an LLM fine-tuned with healthcare data, has performed comparably to human clinical professionals on medical knowledge retrieval, reasoning, and question answering. This suggests that integrating Med-PaLM 2 into medical data annotation workflows could lead to more accurate and efficient annotation processes, harnessing the power of LLMs in the medical domain.
@@
While LLMs have showcased the ability to produce grammatically accurate and human-like text (Thirunavukarasu et al., 2023; Gilson et al., 2022), they still encounter challenges in specialized healthcare contexts. A noticeable knowledge gap in healthcare text still remains between general-domain FMs and medicine professionals, resulting in inaccurate responses to realistic patient inquiries, such as in cardiovascular disease prevention (Sarraju et al., 2023). Besides, Liao et al. (Liao et al., 2023) also highlight that the linguistic attributes of medical text generated by ChatGPT considerably differ from those of human experts.
@@
Advancements are required to improve the utilization of text-based FMs for improving the performance of medical text annotation. PathAsst (Sun et al., 2023) leverages FMs as a generative AI assistant to transform predictive analytics in pathology. Using ChatGPT and GPT-4, PathAsst generates over 180,000 instruction-following samples to invoke pathology-specific models and facilitate effective interactions based on input images and user inputs. Remarkably, generative FMs have shown the potential of text annotation to assist and enhance pathology diagnosis. Also, Liu et al. (Liu et al., 2023e) develop DeID-GPT as a de-identification framework powered by GPT-4 (OpenAI, 2023), designed to detect and eliminate identification information from clinical text records automatically. In contrast to conventional methods, DeID-GPT utilizes in-context learning, which requires no code or procedural changes when applying the solution to different hospitals, different languages, and different data formats. It improves reliability by concealing sensitive information while preserving the original structure of the clinical text. Moreover, Tang et al. (Tang et al., 2023b) utilize the language model BART (Lewis et al., 2019) to enhance medical dialogue generation by focusing on domain-specific terminology. Empirical results highlight the model’s efficiency in generating responses based on the historical context of dialogue exchanges between medical professionals and patients.
@@ 6.2. Medical Image Annotation
Medical image annotation includes outlining and labeling anatomical structures, such as organs, tumors, blood vessels, or bones in histopathology and radiological images (Gao et al., 2021, 2022). Such a task can also involve annotating regions of interest in histopathological slides to identify cancerous cells or specific tissue types of clinical significance. Medical image annotation plays a pivotal role in modern healthcare to help extract and interpret valuable information from complex medical image examinations. This process is essential for both clinical diagnosis and research purposes, as healthcare professionals require computer-aided algorithms to efficiently identify, search, manage, and interpret disease indicators (Yin and Bi, 2018).
Robust segmentation models can be transformed into powerful tools for healthcare image annotation, thus leading to a substantial reduction in annotation costs and the facilitation of large-scale dataset curation spanning a wider range of healthcare imaging applications (Gao et al., 2023). For instance, Qu et al. (Qu et al., 2023) introduce an AI-driven systematic methodology that accelerates organ segmentation annotation by 500 times. While there have been advances in medical image annotation using FMs, the absence of a consensus on state-of-the-art benchmarks remains a significant challenge.
@@
Although SAM is a promising segmentation-focused FM, the inherent domain gap between general vision and medical imaging remains to be addressed. Huang et al. (Huang et al., 2023b) explores various evaluation strategies for SAM. Their analysis delves into the diverse factors affecting SAM’s segmentation performance, unraveling insights crucial for refining its application within the medical context. The outcome consistency and reliability of directly using SAM in medical image segmentation still need to be improved by properly refining FMs (Zhang and Jiao, 2023). Several studies have focused on adapting FMs to the complexities of medical data, aiming to improve the segmentation performance, which potentially benefits the image annotation performance. Zhang et al. (Zhang et al., 2023f) propose SAMAug to involve input augmentation, where prior maps are integrated with raw images to amplify the effectiveness of medical image segmentation. MedSAM exemplifies the process of finetuning SAM on medical images (Ma and Wang, 2023), which ultimately aims to create a versatile tool applicable across a wide range of medical tasks. Similarly, Med SAM Adapter (Wu et al., 2023b) introduces a practice approach by infusing medical-specific domain knowledge into the segmentation model. Such integration significantly enhances the model’s performance within medical image annotation tasks. Extending the range of SAM applications, Poly-SAM (Li et al., 2023b) focuses on polyp segmentation. By fine-tuning the SAM model and comparing transfer learning strategies, this study highlights the substantial potential of adapting SAM to intricate medical image segmentation tasks. Moreover, Hu et al. (Hu et al., 2023b) enhance SAM’s performance through efficient modifications by refining the lightweight task-specific prediction head while keeping the SAM encoder unaltered. Empirical validations indicate that finetuning SAM can significantly elevate its performance on medical image datasets.
Technical challenges remain in adapting SAM to the medical domain efficiently and scalably. Medical images, such as X-rays, MRIs, CT scans, and histopathological slides, could reveal domain-specific clinical patterns that require careful interpretation beyond the scope of general-purpose SAM. In particular, SAM pre-trained in general vision lacks the specialized understanding of medical anatomy, pathology, and radiology necessary for precise image analysis. Data- or parameter-efficient fine-tuning is effective while it is not scalable. Moving forward, more research efforts are required towards a more scalable medical knowledge transfer pattern for better utilizing and adapting SAM or other general-purpose FMs in medical image and text annotation.

### v2

@@ 4. Multi-Modal Data Fusion
Data fusion is a useful strategy to aggregating the information of various medical data modalities towards improved decision-making in healthcare. Standard fusion methods can be conceptually categorized into three types (Stahlschmidt et al., 2022; Huang et al., 2020): early fusion, joint fusion, and late fusion. In early fusion, data from different modalities are combined at the input and passed to the subsequent network (Park et al., 2020; Zhao et al., 2021; Bichindaritz et al., 2021). Joint fusion processes data from each modality through independent networks before fusing their representative feature maps as the input of the subsequent network (Braman et al., 2021; Spasov et al., 2018; Kawahara et al., 2018). Late fusion processes data from each modality through individual networks and fuses their output to make the final outcomes (Reda et al., 2018; Qiu et al., 2018). However, conventional methods suffer from a lack of scalability, generalizability, and cross-modal understanding (Huang et al., 2020). For example, early and joint fusion methods often require sufficient task-specific training data and computational resources, while late fusion methods usually show weaknesses in multi-modal information integration (Budikova et al., 2017). Figure 4 illustrates the healthcare data modalities, data fusion strategies, and associated healthcare tasks.
@@ 4.1. Data Fusing via Joint-Modal Pre-Training
FMs can handle multiple modalities via pre-training on massive-scale paired multi-modal data in a joint-modal mode to obtain a high-level understanding of inter-modality relationships. For instance, CLIP (Radford et al., 2021), ALIGN (Jia et al., 2021), and Florence (Yuan et al., 2021) adopt image-caption pairs collected from the Internet. These models extract image and text features with separated networks and project both features into a shared latent space, enabling a joint training network. These models have shown great zero-shot transferability when testing on unseen image-text combinations, including medical images and healthcare reports (Liu et al., 2023g). In addition, GLIP (Li et al., 2022), CoCa (Yu et al., 2022), and FLAVA (Singh et al., 2022) introduce more advanced fusion modules, such as cross-attention and generative models, to better explore cross-modality interaction during pre-training. Besides vision-language models, there exist models for other modalities with joint-modal pre-training. For example, CLAP (Elizalde et al., 2023) is an audio-language model pre-trained on audio-caption pairs.
Joint-modal pre-training has been applied for healthcare FMs, especially in radiology, where chest X-ray images and reports are frequently paired. ConVIRT (Zhang et al., 2022) is pre-trained on numerous pairs of chest X-rays and radiology reports, fusing vision and language modalities via contrastive learning. BioViL (Boecking et al., 2022) trains a visual model together with an in-domain (CXR) language model by optimizing contrastive loss and masked language modeling. PubMedCLIP (Eslami et al., 2021), CheXzero (Tiu et al., 2022), PLIP (Huang et al., 2023a), and EchoCLIP (Christensen et al., 2024) extend the CLIP model to image-text data in biomedicine, radiology, pathology, and echocardiography respectively. In addition, PRISM (Shaikovski et al., 2024) is trained on thousands of clinical reports, each grouped with one or more whole slide images (WSIs), and can generate reports based on WSIs and text prompts. BiomedJourney (Gu et al., 2023) is a latent diffusion model (Rombach et al., 2022) trained on tuples of pre- and post-disease progression images from patients coupled with a GPT-generated natural language description of the progression, enabling the generation of counterfactual medical images.
@@ 4.2. Data Fusing via LLMs
Transformer-style LLMs possess powerful semantic understanding capability via the attention mechanism (Brown et al., 2020), which can be transferred to multi-modal settings. To be specific, data from different modalities can be aggregated as the prompt input of an LLM (i.e., a sequence of tokens). These combined multi-modal inputs are then fused through the Transformer blocks in the LLM, exchanging information via attention layers (Wang et al., 2023a; Liu et al., 2023a, b). For instance, SimVLM (Wang et al., 2021) fuses image features and text tokens with BERT-like bi-directional attention layers and trains the whole model together. Flamingo (Alayrac et al., 2022) takes as input visual data interleaved with text and inserts gated xattn-dense layers into the language model to condition it on visual inputs. BLIP-2 (Li et al., 2023c) adopts a well-designed Q-Former to bridge the gap between pre-trained vision and language models before passing into the LLM. In addition, LLaVA (Liu et al., 2023b) feeds image features and language instructions into LLaMA to train a multi-modal chatbot. Besides vision and language, recent large multi-modal models (LMMs) are extending to more modalities. For example, NExT-GPT (Wu et al., 2023a) is a unified agent understanding multiple modalities, including text, image, video, and audio. GPT-4o (OpenAI, 2024), which is an optimized version of GPT-4, works with any combination of text, audio, image, and video quickly, enabling more natural human-computer interaction.
@@
Since LLMs show great potential in the medical field (Nori et al., 2023), these methods are applied to medical data fusion as well. Med-Flamingo (Moor et al., 2023b) is an open-ended MedVQA model based on Flamingo, pre-trained on paired and interleaved medical image-text data. MedVInT (Zhang et al., 2023d) applies instruction tuning to a LLaVA-like multi-modal model with MedVQA data to pre-train a generalizable MedVQA foundation model. Visual Med-Alpaca (Shu et al., 2023) is designed to handle multi-modal biomedical tasks, based on the LLaMa-7B architecture and trained with an instruction set designed cooperatively by GPT-3.5-Turbo and human experts. Moreover, LLaVA-Med (Li et al., 2023e) curates biomedical instruction-tuning data by prompting language-only GPT-4 to generate multi-round questions and answers about biomedical images. Surprisingly, GPT-4 generates high-quality visual question-answering conversations even if it has access only to the text. CheXagent (Chen et al., 2024) bridges a vision encoder with a clinical LLM by adopting a Q-Former (Li et al., 2023c), and trains with curated instruction data. These models show outstanding zero-shot/few-shot performance on downstream medical tasks that require multi-modality understanding.
@@
Current multi-modal FMs predominantly focus on language and vision, given their ubiquitous use as the most common modalities. For FM applications in the healthcare field, there are only a handful of examples including molecules (Liu et al., 2023c), genomics (Ding et al., 2023b), electrocardiogram (Chang et al., 2023a), electroencephalogram (Cui et al., 2023a), and tabular data (Wang et al., 2023c). We believe that joint-modal pre-training techniques, combined with LLMs’ logical reasoning capabilities, could benefit healthcare data fusion of modalities beyond the scope of language and vision.
@@ 5. Data Quantity
Towards applying FMs in the healthcare system safely and responsibly, we must address a variety of data quantity hurdles. The trade-off of data quantity lies between the limited information provided by a constrained dataset and the essential information required for training a robust model on the specific healthcare task. Such a disparity can result in inadequate downstream model training, yielding non-robust, inaccurate, and even unreliable model outcomes. However, public patient data records are often scarce due to the rigorous patient privacy protection regulations. In addition, we must deal with the costly real-world dataset curation processes including data collection, cleaning, and annotation. While deep learning has shown its promise, the barrier of limited data access in healthcare remains a significant roadblock for scalable AI-based research.
@@ 5.2. Data Efficiency
General FMs can efficiently retrieve external information from public data sources for downstream tasks using retrieval-augmented generation (RAG) techniques. RAG is a highly effective method for enabling FMs to acquire grounded domain knowledge which is publicly accessible but not included in model pre-training. For instance, BioRAG (Wang et al., 2024) leverages an LLM to adaptively select knowledge sources from papers, search engines, and biological data hubs for biological question-reasoning. REALM (Zhu et al., 2024) extracts comprehensive representations for EHR data by fusing EHR with information retrieved from a medical knowledge graph, without training on task-specific data.
@@ 6.1. Healthcare Text Annotation
Healthcare text annotation is a process that extracts and categorizes critical information within a wide range of healthcare texts. This process plays a crucial role in the healthcare system to enhance data quality and knowledge discovery (Xia and Yetisgen-Yildiz, 2012). However, healthcare text annotation is labor-intensive and time-consuming, observed from studies in medical condition labeling (Briesacher et al., 2008), patient record annotation (Hogan and Wagner, 1997), and key medical entity identification (Abacha and Zweigenbaum, 2011). To streamline a text annotation process, an emerging direction is to harness general-purpose LLMs. For instance, Gilson et al. (Gilson et al., 2022) reveal that ChatGPT (OpenAI, 2021) achieves comparable human-level performance on medical question-answering tasks, showcasing the potential for LLMs to enhance the efficiency of text annotation in a more flexible manner. Med-PaLM 2 (Singhal et al., 2023a), an LLM fine-tuned with healthcare data, has performed comparably to human clinical professionals on medical knowledge retrieval, reasoning, and question answering. This suggests that integrating Med-PaLM 2 into medical data annotation workflows could lead to more accurate and efficient annotation processes, harnessing the power of LLMs in the medical domain.
@@
While LLMs have showcased the ability to produce grammatically accurate and human-like text (Thirunavukarasu et al., 2023; Gilson et al., 2022), they still encounter challenges in specialized healthcare contexts. A noticeable knowledge gap in healthcare text remains between general-domain FMs and medicine professionals, resulting in inaccurate responses to realistic patient inquiries, such as in cardiovascular disease prevention (Sarraju et al., 2023). Besides, Liao et al. (Liao et al., 2023) also highlight that the linguistic attributes of medical text generated by ChatGPT considerably differ from those of human experts.
@@
Advancements are required to improve the utilization of text-based FMs for more accurate and efficient medical text annotation. PathAsst (Sun et al., 2023) leverages FMs as a generative AI assistant to transform predictive analytics in pathology. Using ChatGPT and GPT-4, PathAsst generates over 180,000 instruction-following samples to invoke pathology-specific models and facilitate effective interactions based on input images and user inputs. Remarkably, generative FMs have shown the potential of text annotation to assist and enhance pathology diagnosis. Also, Liu et al. (Liu et al., 2023f) develop DeID-GPT as a de-identification framework powered by GPT-4 (OpenAI, 2023), designed to detect and eliminate identification information from clinical text records automatically. Similarly, Wu et al. (Wu et al., 2024) construct a medical knowledge graph by prompting an LLM to identify and extract entities from segmented medical documents. The LLM is further employed to detect relevance and link the entities, forming a three-level hierarchy. Moreover, Tang et al. (Tang et al., 2023b) utilize a language model to enhance medical dialogue generation by focusing on domain-specific terminology. Empirical results highlight the model’s efficiency in generating responses based on the historical context of dialogue exchanges between medical professionals and patients.
@@ 6.2. Medical Image Annotation
Medical image annotation includes outlining and labeling anatomical structures, such as organs, tumors, blood vessels, or bones in histopathology and radiological images (Gao et al., 2021, 2022). Such a task can also involve annotating regions of interest in histopathological slides to identify cancerous cells or specific tissue types of clinical significance. Medical image annotation is essential in modern healthcare as it helps to extract and interpret valuable information from complex medical image examinations. This is necessary for clinical diagnosis and research, as healthcare professionals require computer-aided algorithms to efficiently identify, search, manage, and interpret disease indicators (Yin and Bi, 2018).
Robust segmentation models can be transformed into powerful tools for healthcare image annotation, leading to a substantial reduction in annotation costs and the facilitation of large-scale dataset curation spanning a wider range of healthcare imaging applications (Gao et al., 2023). For instance, Qu et al. (Qu et al., 2023) introduce an AI-driven systematic methodology that accelerates organ segmentation annotation by 500 times. While there have been advances in medical image annotation using FMs, the absence of a consensus on state-of-the-art benchmarks remains a significant challenge.
@@
Although SAM is a promising segmentation-focused FM, the inherent domain gap between general vision and medical imaging remains to be addressed. Huang et al. (Huang et al., 2023b) explores various evaluation strategies for SAM. Their analysis delves into the diverse factors affecting SAM’s segmentation performance, unraveling insights crucial for refining its application within the medical context. The outcome consistency and reliability of directly using SAM in medical image segmentation still need to be improved by properly refining FMs (Zhang and Jiao, 2023). Several studies have focused on adapting FMs to the complexities of medical data, aiming to improve segmentation performance, which potentially benefits image annotation. Zhang et al. (Zhang et al., 2023f) propose SAMAug to involve input augmentation, where prior maps are integrated with raw images to amplify the effectiveness of medical image segmentation. MedSAM exemplifies the process of finetuning SAM on medical images (Ma and Wang, 2023), which ultimately aims to create a versatile tool applicable across a wide range of medical tasks. Similarly, Med SAM Adapter (Wu et al., 2023b) introduces a practice approach by infusing medical-specific domain knowledge into the segmentation model. Such integration significantly enhances the model’s performance within medical image annotation tasks. Extending the range of SAM applications, Poly-SAM (Li et al., 2023b) focuses on polyp segmentation. By fine-tuning the SAM model and comparing transfer learning strategies, this study highlights the substantial potential of adapting SAM to intricate medical image segmentation tasks. Moreover, Hu et al. (Hu et al., 2023b) enhance SAM’s performance through efficient modifications by refining the lightweight task-specific prediction head while keeping the SAM encoder unaltered. Empirical validations indicate that finetuning SAM can significantly elevate its performance with medical images.
Technical challenges remain in adapting SAM to the medical domain efficiently and scalably. Medical images, such as X-rays, MRIs, CT scans, and histopathological slides, could reveal domain-specific clinical patterns that require careful interpretation beyond the scope of general-purpose SAM. In particular, SAM pre-trained in general vision lacks the specialized understanding of medical anatomy, pathology, and radiology necessary for precise image analysis. Data- or parameter-efficient fine-tuning is effective while not scalable. Moving forward, more research efforts are required towards a more scalable medical knowledge transfer pattern for better utilizing and adapting SAM or other general-purpose FMs in medical image and text annotation.